OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3)

OpenStudy (aroub):

parallel=same slope y=mx+b , m=2 , (5,3) 3=2(5)+b 3=10+b -7=b y=2x-7

OpenStudy (anonymous):

y = 2x + 2 if the equation is parallel to the equation above, then it should be in the form of y = 2x + c by substituting the point (5,3) into the this general equation, c = 3 - 10 = -7 so that the equation is y = 2x - 7. Wish you like this method.

OpenStudy (anonymous):

thank you both! (: